9/10 -1/4- 1/2= ?
anyone can give me the answer and the steps it took to get the answer thanks

Answer by Alan3354(69443)

9/10 -1/4- 1/2= ?
-----------------
Find a common denominator and change all the fractions to that DEN.
It doesn't matter if it's the LCD or not. On a test, don't waste time finding the LCD, just multiply them all together.
------------
10*4*2 = 80
--> 72/80 - 20/80 - 40/80
= 12/80
= 3/20
-------